A consumer electronics company sells 6.7 million MP3 players per month at a price of $100 apiece; when the price is lowered to $80, the company sells 10.2 million players per month. How many MP3 players can the company expect to sell per month at a price of $75?
Answer by stanbon(48568)

A consumer electronics company sells 6.7 million MP3 players per month at a price of $100 apiece; when the price is lowered to $80, the company sells 10.2 million players per month. How many MP3 players can the company expect to sell per month at a price of $75?
----------------------------------
You have two points : (100,6.7) and (80,10.2)
-----------------------
slope = (10.2-6.7)/(80-100) = 3.5/-20 = -7/40
---------------------
equation form: # = (slope)*(price) + b
Substitute to solve for "b":
6.7 = (-7/40)*100 + b
b = 24.20
-------------------
EQUATION:
Let price be "x"; Let # sold be "y"
f(x) = (-7/40)x+24.20
-----------------------
f(75) = (-7/40)*75+24.2
f(75) = 11.0750 million
----------------
Conclusion: When the price is $75 the # sold is 11.075 million
=====================
Cheers,
Stan H.